OpenStudy (anonymous):

I am having trouble with this problem The cost, in millions of dollars, to remove x % of pollution in a lake modeled by c=6000/200-2x What is the cost to remove 75% of the pollutant?

OpenStudy (anonymous):

Plug in x = 75 c = 6000/(200-2(75)) c= 6000/50 c= 120 million dollars! Don't forget to fan!